# Enhancing the Transferability of Adversarial Attacks on Face Recognition with Diverse Parameters Augmentation

### v1

@@ 4.3 Ablation Studies
The hyper-parameter analysis on the $c$ value. The value of $c$ determines the number of ensembles in our proposed attack method, which significantly affects its performance. Hence, we conduct ablation studies on $c$ using the LFW dataset with MobileFace as the surrogate model. To further verify the effectiveness of diverse parameters in enhancing transferability, we select two types of attack methods for comparison. Firstly, we use MobileFace models fine-tuned by a pre-trained backbone and a randomly initialized head in each epoch to generate adversarial examples. We term this adversarial attack method ’Single’. Secondly, we choose the models trained by ‘Single’ and MobileFace models trained by a randomly initialized backbone and head in each epoch to create adversarial examples, implying that the parameters of the trained models are more diverse. We term this attack method ‘Diverse’. The average ASR on IR152, IRSE50, FaceNet, and MobileFace is demonstrated in the left plot of Fig. 7. The left plot of Fig. 7 shows that the ASR increases and then converges as $c$ increases. To analyze the reason, we need to consider the property of $c$. $c$ determines the number of models to be aggregated. If more models are aggregated in each training epoch, the aggregation capacity will increase. If $c$ continues to increase, due to the similarity of the aggregated models in the later epochs, the ASR converges. Moreover, the left plot of Figure 7 demonstrates that the performance of ‘Diverse’ is higher than that of ‘Single’, which verifies the effectiveness of parameter diversity in improving transferability of crafted adversarial examples.
The hyper-parameter analysis on the $\eta$ value. The $\eta$ value is the step size of beneficial perturbations, which is a key hyperparameter in our proposed attack method. We will conduct ablation studies on this parameter using the LFW dataset with MobileFace as the surrogate model. The average ASR on IR152, IRSE50, FaceNet, and MobileFace is shown in the right plot of Fig. 7. To assess the effectiveness of hard models in enhancing the transferability of adversarial examples, we use the Diverse Model Aggregation (DMA) as a baseline for comparison. DMA replaces the hard models in our method with their corresponding vanilla models. From the right plot of Fig. 7, we observe that as the step size of beneficial perturbations increases, the ASR initially rises and then declines. To understand this behavior, we should consider the nature of beneficial perturbations. These perturbations are added to the feature maps of FR models to increase loss when crafting adversarial examples, effectively transforming FR models into hard models. Increasing the step size initially boosts transferability by strengthening the transition to hard models. However, further increasing the step size can degrade the features in the feature maps during forward propagation, ultimately reducing overall attack performance. Additionally, the right plot of Fig. 7 demonstrates that the optimal performance of our proposed method surpasses that of DMA, further validating the effectiveness of the hard model ensemble in our attack method.
Figure 7: The hyper-parameter analysis on the (a) $c$ and (b) $\eta$.
The ablation studies on each stage of our proposed attack method. Our proposed attack method consists of two stages. To verify the effectiveness of each stage, we conduct ablation studies on the stages. Initially, we craft adversarial examples using only the surrogate model, which we denote as ‘Vanilla’. Next, we generated adversarial examples using the ensemble of surrogate models obtained in the DPO stage (i.e. ${\mathbf{F}}$), denoted as DPO. After incorporating the HMA stage, the complete attack method is denoted as ‘DPO + HMA’. The results are presented in Tab. 6. Tab. 6 demonstrates that the addition of the DPO stage results in an increase in black-box ASR, showcasing the effectiveness of the DPO stage in enhancing transferability. Further incorporation of the HMA stage leads to an additional improvement in attack performance, underscoring the effectiveness of the HMA stage in boosting black-box performance. These results collectively demonstrate the effectiveness of each stage in our proposed attack method in improving the transferability.
@@
The ablation studies on the effectiveness of multiple surrogate models at different epochs in improving transferability. Since our proposed attack method utilizes models from intermediate training epochs to craft adversarial examples, it is essential to verify the effectiveness of this approach compared to using only models from the final training epoch. We conduct ablation experiments on these two approaches using the LFW dataset and MobileFace as the surrogate model. We employ the DPO process to obtain the parameter sets ${\mathbf{V}}_{c}^{q}$ and ${\mathbf{V}}_{c}^{m}=\left\{{\mathbf{v}}_{0}^{p},{\mathbf{v}}_{c}^{p},{\mathbf{v}}_{c}^{a}\right\}$. Next, we map ${\mathbf{V}}_{c}^{m}$ into its corresponding model set ${\mathbf{F}}^{m}=\Psi\left({\mathbf{V}}_{c}^{m}\right)$. We then craft adversarial examples by ensembling the models from ${\mathbf{V}}_{c}^{q}$ (i.e. ${\mathbf{F}}$) and ${\mathbf{F}}^{m}$, respectively. The results are presented in Tab. 7. Tab. 7 shows that the performance of the ensemble of ${\mathbf{F}}$ surpasses that of the ensemble of ${\mathbf{F}}^{m}$, demonstrating the effectiveness of using models from intermediate training epochs to enhance the transferability.
@@ 5 Conclusion
We presents a innovative advancement in the field of adversarial attacks on FR through the introduction of the Diverse Parameters Augmentation (DPA) attack method. By addressing the problems of traditional adversarial attacks, particularly the lack of diverse parameters and the exclusion of the FR head, DPA enhances the transferability of adversarial face examples. The extensive experimental results demonstrate the effectiveness of our proposed DPA attack.
@@ 6 Appendix
  • Sec. 6.1. Optimization Objective of Adversarial Attacks in Our Research.
  • Sec. 6.2. Computation Methodology for Attack Success Rate.
  • Sec. 6.3. More Detailed Attack Settings.
  • Sec. 6.4. Comparison Studies on CelebA-HQ.
  • Sec. 6.5. Visual Quality Study.
6.1 Optimization Objective of Adversarial Attacks in Our Research
The objective of adversarial attack as delineated in this paper is as follows:
@@ 6.3 More Detailed Attack Settings
For the tables and figures mentioned—Tab. 1, Tab. 2, Tab. 3, Tab. 4, Tab. 8, Tab. 9, Tab. 10, Tab. 11, Fig. 8, Fig. 5, and Fig. 6, we determine setting $c$ to 35, which corresponds to the optimal value from the left plot in Fig. 7, and setting the step size of beneficial perturbations $\eta$ to 8e-4, as indicated by the optimal value from the right plot in Fig. 7.
Regarding the bottom portion of Fig. 1, we have configured the settings for LGV according to the same hyperparameters as specified in Tab. 5. Similarly, the settings for DI, BPFA, and DPA are aligned with those detailed in Tab. 3.
@@ 6.4 Comparison Studies on CelebA-HQ
To validate the efficacy of our proposed attack method, we create adversarial examples utilizing the CelebA-HQ dataset. The black-box performance of our approach, which employs IRSE50, FaceNet, MobileFace, and IR152 as surrogate models on the CelebA-HQ dataset, is presented in Tab. 8, Tab. 9, Tab. 10, and Tab. 11, respectively. These results consistently indicate that our method outperforms the baseline attacks, thereby highlighting its effectiveness in improving the transferability of adversarial examples.
@@
Table 8: Comparisons of black-box ASR (%) results for attacks using IRSE50 as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
@@
Attacks I F M Iadv Sadv Fadv Madv
FIM 36.3 16.2 80.5 10.5 21.6 5.1 9.5
DI 59.3 42.8 97.6 20.5 39.4 12.0 28.5
DFANet 45.9 25.5 97.0 14.0 30.8 6.6 15.4
VMI 56.7 31.9 96.6 18.5 37.5 9.9 24.3
SSA 58.3 34.8 97.5 19.1 38.2 8.9 22.1
SIA 60.7 40.0 97.9 20.1 40.1 11.5 26.4
BPFA 56.8 27.9 95.3 16.7 32.8 7.5 17.3
BSR 35.7 19.9 86.4 11.2 20.4 5.3 12.0
Ours 68.9 81.7 98.0 40.2 60.6 25.5 53.8
@@
Table 9: Comparisons of black-box ASR (%) results for attacks using FaceNet as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
@@
Attacks I S M Iadv Sadv Fadv Madv
FIM 10.7 16.5 9.6 7.2 8.2 13.0 4.4
DI 22.8 30.4 22.9 15.0 19.9 21.7 11.8
DFANet 15.2 24.3 19.7 10.0 14.1 18.0 7.7
VMI 26.4 36.4 25.7 19.5 25.2 25.3 16.4
SSA 23.5 41.5 36.1 14.9 19.4 22.4 13.2
SIA 29.3 44.5 35.7 21.5 26.0 27.6 18.4
BPFA 20.0 31.0 21.7 12.1 14.0 16.8 7.9
BSR 27.8 43.9 34.0 19.2 25.9 26.0 18.3
Ours 35.5 56.9 58.7 29.9 36.2 32.2 28.3
@@
Table 10: Comparisons of black-box ASR (%) results for attacks using MobileFace as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
@@
Attacks I S F Iadv Sadv Fadv Madv
FIM 7.2 69.0 8.0 3.3 5.4 2.4 12.0
DI 25.0 97.0 32.0 11.0 21.4 8.2 39.0
DFANet 10.7 85.0 12.6 3.7 9.1 2.9 18.8
VMI 20.2 94.7 19.6 7.8 15.8 4.4 31.5
SSA 22.0 95.4 21.0 7.9 15.2 5.3 33.1
SIA 25.4 96.6 25.9 8.9 18.4 5.9 35.8
BPFA 20.7 94.7 17.5 6.7 13.8 4.5 29.7
BSR 10.8 77.1 11.9 3.8 8.1 2.8 15.1
Ours 59.6 97.2 83.5 37.1 57.5 27.4 61.7
@@
Table 11: Comparisons of black-box ASR (%) results for attacks using IR152 as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
Attacks S F M Iadv Sadv Fadv Madv
FIM 39.2 14.1 12.6 18.3 12.1 4.3 4.9
DI 57.6 27.6 27.0 30.1 20.5 8.5 10.8
DFANet 61.2 21.5 22.4 26.6 17.0 5.7 7.9
VMI 56.9 26.4 27.7 32.4 26.1 12.6 16.6
SSA 62.1 23.3 30.7 30.6 19.4 9.1 10.3
SIA 60.8 26.9 30.4 30.4 24.3 12 14.2
BPFA 54.6 15.8 16.4 22.0 14.2 5 5.4
BSR 42.9 17.3 15.0 21.1 15.4 6.6 6.8
Ours 98.0 82.4 95.1 53.5 61.5 27.6 52.8

### v3

@@ 4.3 Ablation Studies
The ablation studies on each stage of our proposed attack method. Our proposed attack method consists of two stages. To verify the effectiveness of each stage, we conduct ablation studies on the stages. Initially, we craft adversarial examples using only the surrogate model, which we denote as ‘Vanilla’. Next, we generate adversarial examples using the ensemble of surrogate models obtained in the DPO stage (i.e. ${\mathbf{F}}$), denoted as DPO. After incorporating the HMA stage, the complete attack method is denoted as ‘DPO + HMA’. The results are presented in Table 3. Table 3 demonstrates that the addition of the DPO stage results in an increase in black-box ASR, showcasing the effectiveness of the DPO stage in enhancing transferability. Further incorporation of the HMA stage leads to an additional improvement in attack performance, underscoring the effectiveness of the HMA stage in boosting black-box performance. These results collectively demonstrate the effectiveness of each stage in our proposed attack method in improving the transferability.
@@
The ablation studies on the effectiveness of multiple surrogate models at different epochs in improving transferability. Since our proposed attack method utilizes models from intermediate training epochs to craft adversarial examples, it is essential to verify the effectiveness of this approach compared to using only models from the final training epoch. We conduct ablation experiments on these two approaches using the LFW dataset and MobileFace as the surrogate model. We employ the DPO process to obtain the parameter sets ${\mathbf{V}}_{c}^{q}$ and ${\mathbf{V}}_{c}^{m}=\left\{{\mathbf{v}}_{0}^{p},{\mathbf{v}}_{c}^{p},{\mathbf{v}}_{c}^{a}\right\}$. Next, we map ${\mathbf{V}}_{c}^{m}$ into its corresponding model set ${\mathbf{F}}^{m}=\Psi\left({\mathbf{V}}_{c}^{m}\right)$. We then craft adversarial examples by ensembling the models from ${\mathbf{V}}_{c}^{q}$ (i.e. ${\mathbf{F}}$) and ${\mathbf{F}}^{m}$, respectively. The results are presented in Table 4. Table 4 shows that the performance of the ensemble of ${\mathbf{F}}$ surpasses that of the ensemble of ${\mathbf{F}}^{m}$, demonstrating the effectiveness of using models from intermediate training epochs to enhance the transferability.
@@
The hyper-parameter analysis studies of our proposed method are in the supplementary.
@@ 5 Conclusion
We present an innovative advancement in the field of adversarial attacks on FR through the introduction of the Diverse Parameters Augmentation (DPA) attack method. By addressing the problems of traditional adversarial attacks, particularly the lack of diverse parameters and the exclusion of the FR head, DPA enhances the transferability of adversarial face examples. In addition, DPA emphasizes the importance of integrating both pre-trained and randomly initialized parameters, providing a fresh perspective on augmentation-based adversarial attacks. The extensive experimental results demonstrate the effectiveness of our proposed DPA attack across multiple datasets and target models. We anticipate that our proposed method will illuminate potential directions for advancing research in adversarial attacks on FR.
6 Acknowledgment
This work was supported in part by the Natural Science Foundation of China under Grant 62372203 and 62302186, in part by the Major Scientific and Technological Project of Shenzhen (202316021), in part by the National key research and development program of China (2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042).
@@ 7.2 More Detailed Attack Settings
For the tables and figures mentioned—Table 6, Table 7, Table 1, Table 5, Table 9, Table 10, Table 11, Table 8, Figure 8, Figure 5, and Figure 6, we determine setting $c$ to 35, which corresponds to the optimal value from the left plot in Figure 7, and setting the step size of beneficial perturbations $\eta$ to 8e-4, as indicated by the optimal value from the right plot in Figure 7.
Regarding the bottom portion of Figure 1, we have configured the settings for LGV according to the same hyperparameters as specified in Table 2. Similarly, the settings for DI, BPFA, and DPA are aligned with those detailed in Table 1.
7.3 More Comparison Studies on LFW
To validate the efficacy of our proposed attack method, we generate adversarial examples using IR152, IRSE50, and FaceNet as surrogate models on the LFW dataset. The black-box performance is presented in Table 5, Table 6, and Table 7, respectively. Our method consistently outperforms baseline attacks, demonstrating its effectiveness in improving the transferability of adversarial examples.
Table 5: Comparisons of black-box ASR (%) results for attacks using IR152 as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
Attacks S F M Iadv Sadv Fadv Madv FIM [71] 29.0 9.3 5.6 13.8 6.8 3.6 2.4 DI [62] 46.9 21.7 14.4 28.0 12.4 7.9 6.1 DFANet [72] 50.7 15.5 12.5 25.6 11.0 5.8 3.2 VMI [53] 49.7 23.9 18.7 30.1 18.3 12.8 11.2 SSA [27] 55.0 21.9 24.0 28.8 14.2 9.0 6.1 SIA [55] 52.6 26.3 19.6 29.8 18.3 11.0 9.5 BPFA [74] 46.7 12.9 9.2 20.1 8.9 4.7 3.1 BSR [46] 35.3 14.7 7.3 19.2 9.9 6.6 4.3 Ours 99.4 90.3 96.4 74.0 69.9 42.0 57.7
Table 6: Comparisons of black-box ASR (%) results for attacks using IRSE50 as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
Attacks I F M Iadv Sadv Fadv Madv FIM [71] 32.3 15.5 79.1 9.8 17.5 5.5 5.7 DI [62] 59.9 47.5 97.7 25.9 41.5 15.6 23.8 DFANet [72] 44.3 26.7 96.9 15.3 28.0 8.6 12.4 VMI [53] 54.0 34.0 96.4 22.5 37.6 13.1 20.3 SSA [27] 58.8 37.1 97.3 22.4 38.5 12.3 18.1 SIA [55] 58.4 41.4 98.2 22.2 37.6 13.9 23.3 BPFA [74] 54.4 27.5 94.6 17.6 29.4 8.1 12.8 BSR [46] 28.7 18.4 84.5 9.2 16.3 6.5 7.6 Ours 74.4 89.8 98.3 57.9 68.2 38.3 59.7
Table 7: Comparisons of black-box ASR (%) results for attacks using FaceNet as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
Attacks I S M Iadv Sadv Fadv Madv FIM [71] 7.8 12.5 5.4 7.5 6.9 17.2 2.5 DI [62] 18.6 32.2 18.5 18.0 15.8 30.2 9.9 DFANet [72] 12.1 22.2 11.7 11.3 10.4 25.1 5.5 VMI [53] 24.4 35.1 20.7 24.4 23.2 36.3 15.2 SSA [27] 21.6 44.8 30.8 17.7 17.0 31.9 10.9 SIA [55] 29.1 42.9 26.2 28.7 23.9 38.3 16.7 BPFA [74] 17.3 31.6 14.7 13.4 13.0 22.6 7.8 BSR [46] 28.6 42.4 25.9 26.2 24.3 34.2 16.0 Ours 42.6 65.0 56.9 47.3 45.4 54.0 31.1
@@ 7.4 Comparison Studies on CelebA-HQ
To further validate the efficacy of our proposed attack method, we create adversarial examples utilizing the CelebA-HQ dataset. The black-box performance of our approach, which employs IR152, IRSE50, FaceNet, and MobileFace as surrogate models on the CelebA-HQ dataset, is presented in Table 8, Table 9, Table 10, and Table 11, respectively. These results consistently indicate that our method outperforms the baseline attacks, thereby highlighting its effectiveness in improving the transferability of adversarial examples.
@@
Table 8: Comparisons of black-box ASR (%) results for attacks using IR152 as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
@@
Attacks S F M Iadv Sadv Fadv Madv FIM [71] 39.2 14.1 12.6 18.3 12.1 4.3 4.9 DI [62] 57.6 27.6 27.0 30.1 20.5 8.5 10.8 DFANet [72] 61.2 21.5 22.4 26.6 17.0 5.7 7.9 VMI [53] 56.9 26.4 27.7 32.4 26.1 12.6 16.6 SSA [27] 62.1 23.3 30.7 30.6 19.4 9.1 10.3 SIA [55] 60.8 26.9 30.4 30.4 24.3 12 14.2 BPFA [74] 54.6 15.8 16.4 22.0 14.2 5.0 5.4 BSR [46] 42.9 17.3 15.0 21.1 15.4 6.6 6.8 Ours 98.0 82.4 95.1 53.5 61.5 27.6 52.8
@@
Table 9: Comparisons of black-box ASR (%) results for attacks using IRSE50 as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
@@
Attacks I F M Iadv Sadv Fadv Madv FIM [71] 36.3 16.2 80.5 10.5 21.6 5.1 9.5 DI [62] 59.3 42.8 97.6 20.5 39.4 12.0 28.5 DFANet [72] 45.9 25.5 97.0 14.0 30.8 6.6 15.4 VMI [53] 56.7 31.9 96.6 18.5 37.5 9.9 24.3 SSA [27] 58.3 34.8 97.5 19.1 38.2 8.9 22.1 SIA [55] 60.7 40.0 97.9 20.1 40.1 11.5 26.4 BPFA [74] 56.8 27.9 95.3 16.7 32.8 7.5 17.3 BSR [46] 35.7 19.9 86.4 11.2 20.4 5.3 12.0 Ours 68.9 81.7 98.0 40.2 60.6 25.5 53.8
@@
Table 10: Comparisons of black-box ASR (%) results for attacks using FaceNet as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
@@
Attacks I S M Iadv Sadv Fadv Madv FIM [71] 10.7 16.5 9.6 7.2 8.2 13.0 4.4 DI [62] 22.8 30.4 22.9 15.0 19.9 21.7 11.8 DFANet [72] 15.2 24.3 19.7 10.0 14.1 18.0 7.7 VMI [53] 26.4 36.4 25.7 19.5 25.2 25.3 16.4 SSA [27] 23.5 41.5 36.1 14.9 19.4 22.4 13.2 SIA [55] 29.3 44.5 35.7 21.5 26.0 27.6 18.4 BPFA [74] 20.0 31.0 21.7 12.1 14.0 16.8 7.9 BSR [46] 27.8 43.9 34.0 19.2 25.9 26.0 18.3 Ours 35.5 56.9 58.7 29.9 36.2 32.2 28.3
Table 11: Comparisons of black-box ASR (%) results for attacks using MobileFace as the surrogate model on the CelebA-HQ dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
Attacks I S F Iadv Sadv Fadv Madv FIM [71] 7.2 69.0 8.0 3.3 5.4 2.4 12.0 DI [62] 25.0 97.0 32.0 11.0 21.4 8.2 39.0 DFANet [72] 10.7 85.0 12.6 3.7 9.1 2.9 18.8 VMI [53] 20.2 94.7 19.6 7.8 15.8 4.4 31.5 SSA [27] 22.0 95.4 21.0 7.9 15.2 5.3 33.1 SIA [55] 25.4 96.6 25.9 8.9 18.4 5.9 35.8 BPFA [74] 20.7 94.7 17.5 6.7 13.8 4.5 29.7 BSR [46] 10.8 77.1 11.9 3.8 8.1 2.8 15.1 Ours 59.6 97.2 83.5 37.1 57.5 27.4 61.7
Figure 7: The hyper-parameter analysis on the (a) $c$ and (b) $\eta$.
7.5 Hyper-parameter Analysis Studies
The hyper-parameter analysis on the $c$ value. The value of $c$ determines the number of ensembles in our proposed attack method, which significantly affects its performance. Hence, we conduct ablation studies on $c$ using the LFW dataset with MobileFace as the surrogate model. To further verify the effectiveness of diverse parameters in enhancing transferability, we select two types of attack methods for comparison. Firstly, we use MobileFace models fine-tuned by a pre-trained backbone and a randomly initialized head in each epoch to generate adversarial examples. We term this adversarial attack method ‘Single’. Secondly, we choose the models trained by ‘Single’ and MobileFace models trained by a randomly initialized backbone and head in each epoch to create adversarial examples, implying that the parameters of the trained models are more diverse. We term this attack method ‘Diverse’. The average ASR on IR152, IRSE50, FaceNet, and MobileFace is demonstrated in the left plot of Figure 7. The left plot of Figure 7 shows that the ASR increases and then converges as $c$ increases. To analyze the reason, we need to consider the property of $c$. $c$ determines the number of models to be aggregated. If more models are aggregated in each training epoch, the aggregation capacity will increase. If $c$ continues to increase, due to the similarity of the aggregated models in the later epochs, the ASR converges. Moreover, the left plot of Figure 7 demonstrates that the performance of ‘Diverse’ is higher than that of ‘Single’, which verifies the effectiveness of parameter diversity in improving transferability of crafted adversarial examples.
The hyper-parameter analysis on the $\eta$ value. The $\eta$ value is the step size of beneficial perturbations, which is a key hyperparameter in our proposed attack method. We will conduct ablation studies on this parameter using the LFW dataset with MobileFace as the surrogate model. The average ASR on IR152, IRSE50, FaceNet, and MobileFace is shown in the right plot of Figure 7. To assess the effectiveness of hard models in enhancing the transferability of adversarial examples, we use the Diverse Model Aggregation (DMA) as a baseline for comparison. DMA replaces the hard models in our method with their corresponding vanilla models. From the right plot of Figure 7, we observe that as the step size of beneficial perturbations increases, the ASR initially rises and then declines. To understand this behavior, we should consider the nature of beneficial perturbations. These perturbations are added to the feature maps of FR models to increase loss when crafting adversarial examples, effectively transforming FR models into hard models. Increasing the step size initially boosts transferability by strengthening the transition to hard models. However, further increasing the step size can degrade the features in the feature maps during forward propagation, ultimately reducing overall attack performance. Additionally, the right plot of Figure 7 demonstrates that the optimal performance of our proposed method surpasses that of DMA, further validating the effectiveness of the hard model ensemble in our attack method.